OpenStudy (anonymous):

Evaluate: lim (sqrt x+2) -3 / x-7 x->7

OpenStudy (anonymous):

use l'hopital and find the limits.....the answer will be, 1/6

OpenStudy (anonymous):

\[\frac{\sqrt{x+2}-3}{x-7}\] \[=\frac{\sqrt{x+2}-3}{x-7}\times \frac{\sqrt{x+2}+3}{\sqrt{x+2}+3}\]

OpenStudy (anonymous):

\[=\frac{x+2-9}{(x-7)(\sqrt{x+2}+3}\] \[=\frac{x-7}{(x-7)(\sqrt{x+2}+3}\] \[=\frac{1}{\sqrt{x+2}+3}\]

OpenStudy (anonymous):

replace x by 7, get \[\frac{1}{\sqrt{7+2}+3}=\frac{1}{3+3}=\frac{1}{6}\]

OpenStudy (anonymous):

thank you, now i know what i did wrong

OpenStudy (anonymous):

it is true that l'hopital will give this as well, but usually these problems come before l'hopital and before derivatives

OpenStudy (anonymous):

yw